# Multi-Layered Reasoning from a Single Viewpoint for Learning See-Through Grasping

### v4

@@ 3.2 Formalizing Multi-Modal VBSeeThruP
Suppose the in-finger camera is used for VBDeformP with an Illuminated Chamber (IC), the pixels of the image are related to the state of the soft membrane captured within the camera’s viewing range, containing zero information about the scene before contact begins [35, 21].
However, for an image taken by an in-finger camera for VBSeeThruP with a See-Thru-Network (STN), the resultant image contains the scene occluded by the Soft Polyhedral Network (SPN). We denote the image as if the SPN is in a Static State before any contact. We can add a hat to use the symbol to indicate a particular case if the scene is empty, captured as uniform pixel features. Here, the subscript timestep is removed to simplify the representation. Since the valuable information captured in this image is purely regarding the SPN, one can add a transparency channel to obtain , represented in white color in Fig. 3. Pixel data of the SPN is preserved in this added data channel, but the pixel data of the empty scene is set to transparent. Such an operation would produce a matrix that can be visualized as a Static Mask of the segmented SPN. Element-wise matrix multiplication of image and this Static Mask will produce masked image . Note that performing the same operation using raw image will produce the same masked image instead. This also explains the SPN’s See-Through feature in the in-finger vision image occlusion process. The above process generally applies to scenarios when the SPN is in a Static State, meaning that no deformation is induced, as there is no contact event. One can use classical image processing techniques to obtain the above results.
However, the above process becomes complicated when contact events occur at . On one side, the SPN enters a Dynamic State, exhibiting whole-body adaptive deformation during physical contact with the objects. On the other side, the See-Through scene also changes during contact, adding to the complexity of segmentation. The resultant image captured by the in-finger camera can be represented as when contact begins. Let us assume that the Dynamic Mask of the segmented SPN, , could be obtained. Then, one should be able to obtain a masked image of the in-finger vision, , by performing the same pixel-wise multiplication of using this Dynamic Mask. Similarly, this would be equivalent to applying the Dynamic Mask over an unobstructed in-finger vision image . The challenge is that such a Dynamic Mask of the SPN will be intricate to obtain using classical methods, especially in cluttered scenes, which usually require advanced segmentation processing and feature engineering.
@@ 3.2.1 Markerless, Real-time, Large-Scale Deformation Tracking
The first problem is to enable markerless real-time segmentation tracking of the STN’s adaptive deformation during contact, which always covers a considerable image area with large-scale spatial deformations. This problem is equivalent to video object segmentation in computer vision. A significant challenge in addressing this problem is to efficiently manage the feature memory model, avoiding memory explosion while minimizing performance decay for long-term predictions. The latest development in a Unified Memory architecture, the XMem model, exhibits state-of-the-art performance on long-video benchmarking datasets [65].
In this study, we experimented with the XMem model for the in-finger video stream, as shown in Fig. 5, achieving a real-time implementation (30 Hz) of the STN’s large-scale deformation mask in a markerless way.
This is realized by using a Static Mask of the STN as the initial tracking target and the in-finger vision image of the STN after contact. Before any contact event, this Static Mask remains the same and can be conveniently obtained manually or using advanced segmentation algorithms such as SAM [19]. The following formal representation addresses the first research problem in Fig. 4.
We evaluated this mask segmentation over a test dataset of 1,000 samples, using the ground truth masks as a reference. The region (J) and boundary (F) measures [66] are 0.975 and 0.997, respectively, showing excellent reliability in tracking the soft finger’s spatial deformation masks in a markerless way. In applications, we found that XMem’s mask tracking of the STN’s deformation is not limited to the Dynamic States but also applicable to the Static States of the STN. We successfully implemented markerless and real-time deformation tracking of the STN whenever our proposed architecture is enabled, alleviating the need for artificial markers reported in previous VBDeformP literature. This process can also be used to detect the occurrence of a contact event, which will be further elaborated upon in the Results section.
@@ 3.2.2 In-Finger Visual Perception from Scene Inpainting
The second problem is reconstructing the scene using occluded images captured by the in-finger vision for visual perception, such as object detection, providing the equivalent capability to an external “Eye” camera. Due to the SPN’s See-Through feature, this problem becomes comparable to inpainting. Video inpainting has recently attracted significant research interest, where learning-based methods exhibit superior performances in filling in missing regions of a video with spatially and temporally coherent contents [67].
@@
We adopted the Flow-Guided Video Inpainting (E2FGVI) algorithm [68], a state-of-the-art algorithm for video inpainting. To facilitate the inpainting process, we obtain a Static Mask of the STN in Fig. 6D using threshold segmentation against a clean background, denoted as . Results show that the STN occludes about 53.6% of the in-finger image pixels. Alternatively, we can interactively obtain the same mask using the Segment Anything Model (SAM) by clicking for comparison in an actual scene [19]. Visual perception is usually performed before the contact event when for an object manipulation pipeline. During this period, the STN is in a Static State, meaning that the mask remains unchanged as the initial mask . The E2FGVI algorithm has been reported with a high Structural Similarity Index (SSI) measure of 0.97 on two public datasets, YouTube-VOS and DAVIS, under stationary masks. Figure 6E shows the scene inpainting pipeline to implement the E2FGVI algorithm using in-finger video clip and the STN’s Static Mask to generate inpainted scene with objects on the tabletop. This inpainting process is represented as
The following focuses on implementing the reactive grasping task in Fig. 1. We use the Real-Time DEtection TRansformer (RT-DETR), a state-of-the-art object detection algorithm [69]. This object detection process is represented as
@@ 3.2.3 Markerless Contact Perception via a Deformation Mask
The third problem is to infer high-performing contact-based perception markerlessly using in-finger vision. This representation problem is usually resolved by adding artificial markers to the soft medium, which VBSeeThruP shares using either an Illuminated Chamber or a See-Thru-Network. Previous research has also explored fiducial markers within the STN, achieving high-performing proprioceptive perception, particularly when considering the STN’s static and dynamic viscoelastic properties [18].
This study utilizes the STN masks in the Dynamic States, obtained from segmented tracking, as the model input to infer 6D forces and torques. We adopted the Supervised Variational AutoEncoder (SVAE) for this task, and Fig. 7A shows the experimental setup for collecting true labels of 6D forces and torques at the finger base using nano25 by ATI for calibration. The in-finger camera and FT sensor were connected to a laptop for data streaming. We collected 40,000 synchronized pairs of images and FT readings while manually compressing the two STN fingers from various directions and contact locations. We used an 8:2 train-valid split and trained the SVAE model for 50 epochs. To avoid excessive good validation scores due to the potential sample dependence, we separately collected a test dataset of 1,000 samples.
@@
The overall loss comprises reconstruction loss, Mean-Squared Error (MSE) loss for force and torque prediction, and Kullback-Leibler divergence. The average time for a single inference is 1.9 milliseconds on an NVIDIA GeForce RTX 3080 Ti Laptop GPU. The Mean Absolute Errors of the validation and test datasets are reported in Table 2, indicating good generalization of the learned model.
Figure 7C compares the test dataset’s prediction and ground truth in time series. Figure 7D shows the scatter plot of such comparison in the validation dataset with excellent scores over 0.985 in all components except . The relatively weak perforce in is due to the limited adaptability along the -axis. We also evaluate the resultant force in the plane and find that the mean magnitude and orientation error are 0.286 N and 2.6 degrees, respectively.
@@ 3.3 Ablation Study
This section presents an ablation study demonstrating how VBSeeThruP contributes to a multi-modal grasping pipeline without requiring additional or external devices. In particular, we investigate how in-finger vision and soft compliance from VBSeeThruP improve the reliability and generalization in a pick-and-place scenario. Please refer to Supplementary Video S2 for further details.
@@
For the calibration process, we place four oranges (or any objects of similar size) on the table and detect their centers in the image plane using YOLO11 [70] (see Fig. 8A). An operator then manually bends the soft finger (for horizontal - movement) to position the gripper over each orange. Once the gripper is centered above an orange, we record the corresponding world coordinates in the robot base frame by twisting the soft finger (clockwise and counter-clockwise to mark the start and end of data recording). The overall setup is shown in Figs. 8B&C, and a full demonstration is provided in the supplementary video. Notably, our method solely relies on VBSeeThruP for depth detection, whereas a typical solution would require an external depth camera.
@@ 3.3.1 Abalation Experiment Setup
SPN + In-finger Vision: We augment the SPN with soft fingers’ in-finger vision for scene inpainting and YOLO11-based [70] object localization. The grasp width remains fixed at .
@@
3.3.2 Abalation Results and Discussion
Each experiment scenario was repeated three times. We observed that even in the most straightforward configuration (SPN Only), objects could often be picked without damaging them or triggering safety stops, owing to the compliance offered by the soft finger. For example, the soft, deformable fingertips are easily adapted to objects from different angles, as shown in Fig. 8E (grasping an orange from the side with bending and twisting deformations). In contrast, using the original rigid fingertips without carefully specifying object dimensions frequently caused collisions and emergency stops.
When in-finger vision was added (SPN + In-finger Vision), the grasp location and orientation were derived from YOLO11, significantly improving the success rate over random grasps. Nevertheless, the lack of force feedback made the system less robust under disturbance, causing the robot to fail regrasp attempts in several trials (shown in the supplementary video). In one case, the fingertip was misaligned and exerted excessive force on a plastic ball, causing it to slide out of the gripper.
The configuration with SPN + In-finger Vision + Inferred 6D FT overcame these issues using 6D force/torque estimates for closed-loop grasping. As a result, it achieved a 100% success rate under the same disturbance conditions. It also enabled weight classification to differentiate between empty balls and allowed the system to adaptively regrasp objects if an object fell was detected.
@@ 4.1 Learning Reactive See-Through Grasping
The task begins with visual perception of the scene using the VBSeeThruP for object classification and localization. During the first attempt, the gripper will attempt to pick up the object at the center of the bounding box and a fixed rotation angle and depth perpendicular to the table. We intentionally used a simpler algorithm without the optimal orientation to test reactive grasping.
While approaching the initial grasping pose, VBSeeThruP detects if it has contact with the object by inferring real-time 6D forces and torques based on the finger’s whole-body deformation. Due to the network design, the finger can deform in a twisted manner to generate torque about the -axis. When is detected to be larger than 0.05 Nm, the object orientation is severely different from the gripper’s initial pose, suggesting a less ideal grasping that can be optimized through regrasping by turning the wrist joint reversely to reduce for an enhanced adaptation and grasping robustness. When is within the threshold, the gripper closes its fingers until the gripping force reaches the predefined maximum gripping force of 3 N. Then, the gripper lifts the object and drops it into the placement box. We repeated the process until all five objects were cleared from the table.
@@
4.2 Towards a Multi-Modal Soft Touch vis VBSeeThruP
Here, we discuss the extended modalities perceivable via our proposed architecture in Fig. 1, providing an intuitive understanding of the three research problems illustrated in Fig. 4 that aim to achieve multi-modal perception with a soft touch. Please refer to the following sections and Supplementary Video S4 for further details.
Using a single sensory input from the in-finger vision featuring See-Thru-Networks, we present a multi-modal approach towards robotic perceptual embodiment via VBSeeThruP with a soft touch, as shown in Fig. 10. Our proposed architecture can be generalized to further generate various outputs for extrinsic perception, including inpainted scenes, object detection, depth sensing, and scene segmentation. For intrinsic perceptions, we demonstrate the implementation of our architecture, which generates outputs including masked deformation tracking, 6D FT sensing in real-time with high accuracy, and contact event detection. Further expansion to other intrinsic or extrinsic modalities can also be implemented.
@@ 4.2.1 Markerless, Real-Time & Large-Scale Deformation Tracking
VBSeeThruP generically describes a class of soft robotic fingers/fingertips that involves a See-Thru-Network with in-finger vision as a design feature. The See-Through feature enables the in-finger vision to simultaneously capture visual features from the finger structure and the external environment. In this study, the See-Thru-Network is implemented using the Soft Polyhedral Network. However, researchers have a wide range of design choices to customize the soft interface as needed, as long as the resultant image contains sufficient details from both the finger structure and the external environment. The resultant performance is determined by the level of detail when extracting the soft structure’s deformation and the reusable information left to reconstruct the scene.
@@
Algorithmic development in video object segmentation is a well-researched field that enables convenient implementation of this capability for See-Thru-Networks with in-finger vision. The 30 Hz mask tracking implemented in this study is sufficiently high to enable real-time processing for robotics without adding to the computational burden of the robotic systems. Future research to allow a higher mask tracking bandwidth is preferred to increase the upper bound in collecting sensory feedback (towards 1 kHz) from in-finger vision, requiring higher-framerate miniature cameras and high-performance mask tracking algorithms for more challenging robotic applications. The quality of mask segmentation at this step generally determines the maximum performance one can expect from the VBSeeThruP architecture in multi-modal perception. The solution to this problem provides the foundation for implementing suitable robotic solutions to address the second and third research problems.
4.2.2 Multi-Modal Visual Perception of the Occluded Scene via In-Finger Vision

### v5

@@ 3.2 Formalizing Multimodal VBSeeThruP
Suppose the in-finger camera is used for VBDeformP with an Illuminated Chamber (IC), the pixels of the image are related to the state of the soft membrane captured within the camera’s viewing range, containing zero information about the scene before contact begins (Yuan et al., 2017a; Li et al., 2024a).
However, for an image taken by an in-finger camera for VBSeeThruP with a See-Thru-Network (STN), the resultant image contains the scene occluded by the Soft Polyhedral Network (SPN). We denote the image as if the SPN is in a Static State before any contact. We can add a hat to use the symbol to indicate a particular case if the scene is empty, captured as uniform pixel features. Here, the subscript timestep is removed to simplify the representation. Since the valuable information captured in this image is purely regarding the SPN, one can add a transparency channel to obtain , represented in white color in Fig. 3. Pixel data of the SPN is preserved in this added data channel, but the pixel data of the empty scene is set to transparent. Such an operation would produce a matrix that can be visualized as a Static Mask of the segmented SPN. Element-wise matrix multiplication of image and this Static Mask will produce masked image . Note that performing the same operation using raw image will produce the same masked image instead. This also explains the SPN’s See-Through feature in the occlusion process for in-finger vision images. The above process generally applies when the SPN is in a Static State, meaning no deformation is induced because there is no contact event. One can use classical image processing techniques to obtain the above results.
However, the above process becomes complicated when contact events occur at . On one side, the SPN enters a Dynamic State, exhibiting whole-body adaptive deformation during physical contact with the objects. On the other hand, the See-Through scene also changes during contact, further complicating segmentation. The resultant image captured by the in-finger camera can be represented as when contact begins. Let us assume that the Dynamic Mask of the segmented SPN, , could be obtained. Then, one should be able to obtain a masked image of the in-finger vision, , by performing the same pixel-wise multiplication of using this Dynamic Mask. Similarly, this would be equivalent to applying the Dynamic Mask over an unobstructed in-finger vision image . The challenge is that a Dynamic Mask of the SPN is intricate to obtain using classical methods, especially in cluttered scenes, which often require advanced segmentation and feature engineering.
@@ 3.2.1 Markerless, Real-time, Large-Scale Deformation Tracking
The first problem is to enable markerless, real-time segmentation and tracking of the STN’s adaptive deformation during contact, which often spans a large image area and involves large-scale spatial deformations. This problem is equivalent to video object segmentation in computer vision. A significant challenge in addressing this problem is efficiently managing the feature memory model, avoiding memory explosion while minimizing performance degradation for long-term predictions. The latest development in a Unified Memory architecture, the XMem model, exhibits state-of-the-art performance on long-video benchmarking datasets (Cheng and Schwing, 2022).
In this study, we experimented with the XMem model for the in-finger video stream, as shown in Fig. 5, achieving real-time (30 Hz) markerless implementation of the STN’s large-scale deformation mask.
This is realized by using a Static Mask of the STN as the initial tracking target and the in-finger vision image of the STN after contact. Before any contact event, this Static Mask remains the same and can be conveniently obtained manually or using advanced segmentation algorithms such as SAM (Kirillov et al., 2023). The following formal representation addresses the first research problem in Fig. 4.
We evaluated this mask segmentation on a test dataset of 1,000 samples, using the ground-truth masks as a reference. The region (J) and boundary (F) measures (Perazzi et al., 2016) are 0.975 and 0.997, respectively, showing excellent reliability in tracking the soft finger’s spatial deformation masks in a markerless way. In applications, we found that XMem’s mask-tracking of the STN’s deformation is not limited to Dynamic States but also applies to Static States. We successfully implemented markerless and real-time deformation tracking of the STN whenever our proposed architecture is enabled, alleviating the need for artificial markers reported in previous VBDeformP literature. This process can also be used to detect the occurrence of a contact event, as further elaborated in the Results section.
@@ 3.2.2 In-Finger Visual Perception from Scene Inpainting
The second problem is reconstructing the scene using occluded images captured by the in-finger vision for visual perception, such as object detection, providing the equivalent capability to an external “Eye” camera. Due to the SPN’s See-Through feature, this problem becomes comparable to inpainting. Video inpainting has recently attracted significant research interest, where learning-based methods exhibit superior performances in filling in missing regions of a video with spatially and temporally coherent contents (Quan et al., 2024).
@@
We adopted the Flow-Guided Video Inpainting (E2FGVI) algorithm (Li et al., 2022), a state-of-the-art algorithm for video inpainting. To facilitate the inpainting process, we obtain a Static Mask of the STN in Fig. 6D using threshold segmentation against a clean background, denoted as . Results show that the STN occludes about 53.6% of the in-finger image pixels. Alternatively, we can interactively obtain the same mask using the Segment Anything Model (SAM) by clicking for comparison in an actual scene (Kirillov et al., 2023). Visual perception is usually performed before the contact event when for an object manipulation pipeline. During this period, the STN is in a Static State, meaning that the mask remains unchanged as the initial mask . The E2FGVI algorithm has been reported to achieve a high Structural Similarity Index (SSI) of 0.97 on two public datasets, YouTube-VOS and DAVIS, under stationary masks. Figure 6E shows the scene inpainting pipeline to implement the E2FGVI algorithm using in-finger video clip and the STN’s Static Mask to generate inpainted scene with objects on the tabletop. This inpainting process is represented as
The following focuses on implementing the reactive grasping task in Fig. 1. We use the Real-Time DEtection TRansformer (RT-DETR), a state-of-the-art object detection algorithm (Zhao et al., 2024). This object detection process is represented as
@@ 3.2.3 Markerless Contact Perception via a Deformation Mask
The third problem is to infer high-performing contact-based perception markerlessly using in-finger vision. This representation problem is usually resolved by adding artificial markers to the soft medium, which VBSeeThruP shares using either an Illuminated Chamber or a See-Thru-Network. Previous research has also explored fiducial markers within the STN, achieving high-performing proprioceptive perception, particularly when considering the STN’s static and dynamic viscoelastic properties (Liu et al., 2024b).
This study uses the STN masks in the Dynamic States, obtained from segmented tracking, as model inputs to infer 6D forces and torques. We adopted the Supervised Variational AutoEncoder (SVAE) for this task, and Fig. 7A shows the experimental setup for collecting true labels of 6D forces and torques at the finger base using nano25 by ATI for calibration. The in-finger camera and FT sensor were connected to a laptop for data streaming. We collected 40,000 synchronized pairs of images and FT readings as we manually compressed the two STN fingers from various directions and contact locations. We used an 8:2 train-valid split and trained the SVAE model for 50 epochs. To avoid excessive validation scores due to potential sample dependence, we collected a separate test dataset of 1,000 samples.
@@
The overall loss comprises reconstruction loss, Mean-Squared Error (MSE) loss for force and torque prediction, and Kullback-Leibler divergence. The average time per inference is 1.9 milliseconds on an NVIDIA GeForce RTX 3080 Ti Laptop GPU. The Mean Absolute Errors of the validation and test datasets are reported in Table 2, indicating good generalization of the learned model.
Figure 7C compares the test dataset’s prediction and ground truth in time series. Figure 7D shows the scatter plot of this comparison in the validation dataset, with excellent scores above 0.985 across all components except . The relatively weak perforce in is due to the limited adaptability along the -axis. We also evaluate the resultant force in the plane and find that the mean magnitude and orientation error are 0.286 N and 2.6 degrees, respectively.
@@ 3.3 Ablation Study
This section presents an ablation study demonstrating how VBSeeThruP contributes to a multimodal grasping pipeline without requiring additional or external devices. In particular, we investigate how in-finger vision and soft compliance from VBSeeThruP improve the reliability and generalization in a pick-and-place scenario. Please refer to Supplementary Video S2 for further details.
@@
For the calibration process, we place four oranges (or any objects of similar size) on the table and detect their centers in the image plane using YOLO11 (Jocher et al., 2024) (see Fig. 8A). An operator then manually bends the soft finger (for horizontal - movement) to position the gripper over each orange. Once the gripper is centered above an orange, we record the corresponding world coordinates in the robot base frame by twisting the soft finger (clockwise and counter-clockwise to mark the start and end of data recording). The overall setup is shown in Figs. 8B&C, and a full demonstration is provided in the supplementary video. Notably, our method solely relies on VBSeeThruP for depth detection, whereas a typical solution would require an external depth camera.
@@ 3.3.1 Experiment Setup
SPN + In-finger Vision: We augment the SPN with soft fingers’ in-finger vision for scene inpainting and YOLO11-based (Jocher et al., 2024) object localization. The grasp width remains fixed at .
@@
3.3.2 Results and Discussion
Each experiment scenario was repeated three times. We observed that even in the most straightforward configuration (SPN Only), objects could often be picked up without damaging them or triggering safety stops, thanks to the soft finger’s compliance. For example, the soft, deformable fingertips are easily adapted to objects from different angles, as shown in Fig. 8E (grasping an orange from the side with bending and twisting deformations). In contrast, using the original rigid fingertips without carefully specifying object dimensions often resulted in collisions and emergency stops.
When in-finger vision was added (SPN + In-finger Vision), the grasp location and orientation were derived from YOLO11, significantly improving the success rate over random grasps. Nevertheless, the lack of force feedback made the system less robust to disturbances, causing the robot to fail regrasp attempts in several trials (as shown in the supplementary video). In one case, the fingertip was misaligned and exerted excessive force on a plastic ball, causing it to slide out of the gripper.
The configuration with SPN + In-finger Vision + Inferred 6D FT overcame these issues using 6D force/torque estimates for closed-loop grasping. As a result, it achieved a 100% success rate under the same disturbance conditions. It also enabled weight classification to differentiate between empty balls and allowed the system to adaptively regrasp objects when an object that had fallen was detected.
@@ 4.1 Learning Reactive See-Through Grasping
The task begins with visual perception of the scene using the VBSeeThruP for object classification and localization. During the first attempt, the gripper will pick up the object at the center of the bounding box, with a fixed rotation angle and a depth perpendicular to the table. We intentionally used a simpler algorithm without the optimal orientation to test reactive grasping.
While approaching the initial grasping pose, VBSeeThruP detects if it has contact with the object by inferring real-time 6D forces and torques based on the finger’s whole-body deformation. Due to the network design, the finger can deform in a twisted manner, generating torque about the -axis. When is detected to be larger than 0.05 Nm, the object orientation is severely different from the gripper’s initial pose, suggesting a less ideal grasping that can be optimized through regrasping by turning the wrist joint reversely to reduce for an enhanced adaptation and grasping robustness. When is within the threshold, the gripper closes its fingers until the gripping force reaches the predefined maximum gripping force of 3 N. Then, the gripper lifts the object and drops it into the placement box. We repeated the process until all five objects were cleared from the table.
@@
4.2 Towards a Multimodal Soft Touch vis VBSeeThruP
Here, we discuss the extended modalities perceivable via our proposed architecture in Fig. 1, providing an intuitive understanding of the three research problems illustrated in Fig. 4 that aim to achieve multimodal perception with a soft touch. Please refer to the following sections and Supplementary Video S4 for further details.
Using a single sensory input from the in-finger vision, featuring See-Thru-Networks, we present a multimodal approach to robotic perceptual embodiment via VBSeeThruP with soft touch, as shown in Fig. 10. Our proposed architecture can be generalized to further generate various outputs for extrinsic perception, including inpainted scenes, object detection, depth sensing, and scene segmentation. For intrinsic perceptions, we demonstrate the implementation of our architecture, which generates outputs including masked deformation tracking, real-time, high-accuracy 6D FT sensing, and contact event detection. Further expansion to other intrinsic or extrinsic modalities can also be implemented.
@@ 4.2.1 Markerless, Real-Time & Large-Scale Deformation Tracking
VBSeeThruP generically describes a class of soft robotic fingers/fingertips that involves a See-Thru-Network with in-finger vision as a design feature. The See-Through feature enables the in-finger vision to simultaneously capture visual features from the finger structure and the external environment. In this study, the See-Thru-Network is implemented using the Soft Polyhedral Network. However, researchers have a wide range of design choices to customize the soft interface as needed, provided the resultant image contains sufficient details from both the finger structure and the external environment. The resultant performance is determined by the level of detail when extracting the soft structure’s deformation and the reusable information left to reconstruct the scene.
@@
Algorithmic development in video object segmentation is a well-researched field that enables convenient implementation of this capability for See-Thru-Networks with in-finger vision. The 30 Hz mask tracking implemented in this study is sufficiently high to enable real-time processing for robotics without adding to the computational burden of the robotic systems. Future research to enable higher mask-tracking bandwidth is preferred to increase the upper bound for collecting sensory feedback (towards 1 kHz) from in-finger vision, requiring higher-framerate miniature cameras and high-performance mask-tracking algorithms for more challenging robotic applications. The quality of mask segmentation at this step generally determines the maximum performance one can expect from the VBSeeThruP architecture in multimodal perception. The solution to this problem provides the foundation for implementing suitable robotic solutions to address the second and third research problems.
4.2.2 Multimodal Visual Perception of the Occluded Scene via In-Finger Vision